OpenStudy (anonymous):

Find a vector perpendicular to Oy axis. Knowing that v*v1=8 and v*v2=-3. v1=(3,1,-2) and v2=(-1,1,1)

OpenStudy (anonymous):

A bit confused by this question, is this what you have written above the dot product? the x-axis is perpendicular to the y-axis, also the z-axis

OpenStudy (turingtest):

...or any other vector in the without a y component

OpenStudy (turingtest):

oh wait, it's supposed to be perpendicular to which axis? @viniterranova please explain so we can help you

OpenStudy (anonymous):

Determine the vector v, orthogonal to the axis Oy.

OpenStudy (anonymous):

where v*v1=8 and v*v2=-3

OpenStudy (turingtest):

you did not really clarify you just repeated yourself are these dot products?

OpenStudy (anonymous):

Yes these are dot products

OpenStudy (anonymous):

then the y component can be left away, will result in a system of equations.

OpenStudy (anonymous):

Here it´s a shot from the book where i taken the question.

OpenStudy (turingtest):

so we have\[\vec v\cdot\vec v_1=8\]\[\vec v\cdot\vec v_2=-3\]where\[\vec v_1=\langle3,1,-2\rangle\]\[\vec v_2=\langle-1,1,1\rangle\]and if it is perpendicular to the y axis we have that\[\vec v\cdot\langle0,1,0\rangle=0\]let\[\vec v=\langle x,y,z\rangle\] and we get the system of equations spacelimbus was referring to

OpenStudy (anonymous):

And, what´s more?

OpenStudy (turingtest):

\[\vec v\cdot\vec v_1=3x+y-2z=8\]\[\vec v\cdot\vec v_2=-x+y+z=-3\]\[\vec v\cdot\langle0,1,0\rangle=y=0\]

OpenStudy (anonymous):

So, in this case y = zero? That ´s right? So, i just plug 0 in the equation in order to slove the system? Am i right?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

In this i got a following system 3x-2z=8 and -x+z= -3 and solving in function of z i got z=x-3 and so, i have reached x=2. Am i right?

OpenStudy (anonymous):

x=2, z=-1, is what I got if you want to compare.

OpenStudy (anonymous):

Yes, that´s for sure. x=2 and z=-1. Thanks guys.

OpenStudy (anonymous):

So the answer to the question is w=(2,0,-1)

OpenStudy (anonymous):

well in the exercise they called the vector v, so you maybe want to name him like that too, but that's not a convention, the coordinates are what's important, and you maybe want to write it in brackets, see how @TuringTest wrote the vectors.

OpenStudy (anonymous):

Thanks for the help. But, i don´t know how to use the Latex yet. Thanks a lot for the help.